OpenStudy (anonymous):

-5n+1-2=-6n+2...how do you do this with an explaination?

OpenStudy (anonymous):

-5n + 1-2 = -6n + 2 -5n - 1 = -6n + 2 We move over the 2 -5n - 1 - 2 = -6n Now we move over the -5n -3 = -6n + 5n -3 = -1n Divide both sides by -1 -3 / -1 = -1n / -1 3 = n n = 3 You should check out this site: http://www.wolframalpha.com/input/?i=-5n%2B1-2%3D-6n%2B2+ :)

OpenStudy (anonymous):

does n = 3? thats what i got for the answer but im not sure

OpenStudy (anonymous):

Thats right!

OpenStudy (anonymous):

haha thankyou so much!(: